a) The arithmetic mean of 10 numbers is 4, when an eleventh number, x is added so that the overall mean is changed to 5. When a twelfth number, y is added the mean changes to 4. Determine the values of x and y
1
Expert's answer
2019-03-28T11:05:47-0400

(4*10+x)/11=5 -> x=5*11-4*10=15.

(5*11+x)/12=4 -> x=4*12-5*11=--7.